PRESENT VALUE OF A PERPETUITY What is the present value of a $700 perpetuity if the interest rate is 10%? Round your answer to the nearest cent. $ If interest rates doubled to 20%, what would its present value be? Round your answer to the nearest cent.
Answer #1

Present value of perpetuity=Annual cash flows/interest rate

1.Present value=700/0.1=$7000

2.Present value=700/0.2=$3500